Solve and Round to the nearest millionth?
log[4]x+log[4](x-5)=0

Answer by stanbon(75887)

Solve and Round to the nearest millionth?
log[4]x+log[4](x-5)=0
----
log4(x(x-5)) = 0
---
x^2-5x-1 = 0
x = [5 +- sqrt(25-4*5*-1)]/2
---
Positive soltion:
x = [5 + sqrt(45)]/2
---
x = [5 + 3sqrt(5)]/2
-----
x = 5.854102
=================
Cheers,
Stan H.
=================